OpenStudy (anonymous):

What is the sum 1/2b + b/2

OpenStudy (anonymous):

it is b u want to know how?

OpenStudy (anonymous):

no its like a fracton. I dont really get how to do them. Do you cross multiply?

OpenStudy (anonymous):

half of book plus another half will give us one book

OpenStudy (anonymous):

First you do common deninator.

OpenStudy (anonymous):

Okay. wouldnt that be 2?

OpenStudy (anonymous):

1/2b+b/2 Find the common denominator, 2b, then 1/2b+ b(b)/2(b). 1/2b+b^2/2b= (b^2+1)/2b Therefore, 1/2b +b/2= (b^2+1)/2b